Webbie10000:

A right rectangular prism is 5and 1/4 in. wide, 12and 1/2 in. long, and 4 in. tall. What is the volume of the prism? 50 in3 65and 5/8 in3 260 in3 262and 1/2 in3

tigerlover:

Volume of rectangular prism = 1 1/3 x 5/6 x 2/3 = 4/3 x 5/6 x 2/3 = 20/27 Volume of cube = 1/6 x 1/6 x 1/6 = 1/216 Number of cubes that will pack the rectangular prism = 20/27 / 1/216 = 160

Webbie10000:

answer is 160

tigerlover:

Sorry. I got confused. Hold on.

Webbie10000:

ok

tigerlover:

Volume = 514 * 1212 * 4 Volume = 2491872 cube inches

Webbie10000:

thx but is that answer choice

tigerlover:

Oh. Hold on. Let me convert the number to inches.

Webbie10000:

ok thx

tigerlover:

I'm so sorry for all the confusion. Please stand by.

Webbie10000:

ok

tigerlover:

Hold on. I figured it out. If we take \[5 \frac{ 1 }{ 4 }* 12 \frac{ 1 }{ 2 } * 4=?\] What does this equal?

Webbie10000:



tigerlover:

I didn't see that there were fractions. Sorry for all the false answers.

Webbie10000:

it ok you can help me now

tigerlover:

This would be \[5 \frac{ 1 }{ 4 }*12\frac{ 1 }{ 2 }*4= 262 \frac{ 1 }{ 2 }\]

Webbie10000:

thx

tigerlover:

No problem! :)

Webbie10000:

tysm i got an B 80%

Webbie10000:

go to my new question thing

tigerlover:

Where?